Assume that the playbook contains 10 passing plays and 13 running plays. The coach randomly selects 8 plays from the playbook.

What is the probability that the coach selects at least 3 passing plays and at least 2 running plays?
Answer #1

probability that the coach selects at least 3 passing plays and at least 2 running plays

=1-P(less than 3 passing or less than 2 running plays)

=1-P(less than 3+less than 2 running plays) =1-(0.195114+0.03273)=0.801613